OpenStudy (anonymous):

a coin jar contains only nickels and quarters. there are exactly 38 coins in the jar, and the total value of the coins is $2.50. what is x and what is y?

OpenStudy (anonymous):

somebody help me here i have a test in 2 days

OpenStudy (anonymous):

We have two equations to solve here, the first is x+y=38, the other is .05x (for the nickels) + .25y (for the quarters) = $2.50. we can find x or y by using the first equation, which I solved for x, so x=38-y. We can use x here for the second equation and we wind up with, \[.05(38-y)+.25y = 2.5\]. Now all you have to do is solve for y and take the difference from the first equation to get the amount of quarters. I got 3 quarters and 35 nickels.